There are currently issues with accessing /fs/ess/ locations using ondemand.

Known issues

Unresolved known issues

Known issue with an Unresolved Resolution state is an active problem under investigation; a temporary workaround may be available.

Resolved known issues

A known issue with a Resolved (workaround) Resolution state is an ongoing problem; a permanent workaround is available which may include using different software or hardware.

A known issue with Resolved Resolution state has been corrected.

Known Issues

Title Category Resolution Description Posted Updated
Rolling reboot of Owens cluster, starting from 9AM September 11, 2017 Batch, Owens Resolved

Updates on 12:20PM September 25, 2017: 

The rolling reboot of Owens is completed. 

... Read more
4 years 11 months ago 4 years 10 months ago
Some issues remain after downtime Login Problems, Operations, Outage Resolved

15 July 2016, 5:00PM update: some additional issues we are facing

  • We are experiencing periodic hangs of the GPFS client file system software used with the new storage environment. We... Read more
6 years 1 month ago 4 years 10 months ago
OnDemand has NOT been working with external providers since 08/22 OnDemand Resolved

Updates on 9:40AM August 23, 2017: this issue has been resolved. 

>>>

Issue:

User can NOT login to OnDemand (ondemand.osc.edu)... Read more

4 years 11 months ago 4 years 11 months ago
PBS commands on Owens are not working Batch, Owens Resolved

Update posted on July 12, 2017 at 1:50PM:

We have fixed the problem with the batch management system on Owens and queues on Owens have been opened again for jobs.

... Read more

5 years 1 month ago 5 years 1 month ago
Rolling reboot of Owens cluster, starting from 9AM June 28, 2017 Owens Resolved

Update posted on July 7, 2017 at 2:00PM:

Rolling reboot of login and compute nodes of Owens cluster is completed. 

... Read more
5 years 1 month ago 5 years 1 month ago
Systemic Problem on Cluster Computing service Operations Resolved

4:20PM 6/23/2017 Update: All HPC systems are back in production. This outage may cause failures of users' jobs. We'll update the community as more is known. 

... Read more
5 years 1 month ago 5 years 1 month ago
my.osc.edu is NOT available Account Management Resolved

my.osc.edu has not been fully restored after yesterday's downtime. You can change your password, but you will not be able to use the new password on my.osc.edu. The updated password will work to... Read more

5 years 2 months ago 5 years 2 months ago
"pbsdcp" is not working on Oakley Resolved

12:35PM 5/24/2017 Update: pbsdcp   has been fixed on Oakley.

pbsdcp   is not working on Oakley and returns a missing library error as below:... Read more

5 years 2 months ago 5 years 2 months ago
Issue with GPFS on Owens since April 14, 2017 Batch, filesystem, Owens Resolved

3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address this GPFS issue. 

We have had issues with GPFS mounts on Owens Cluster since Friday afternoon,... Read more

5 years 3 months ago 5 years 3 months ago
Rolling reboot of all clusters, starting from Wednesday morning, April 19, 2017 Batch, Maintenance, Owens, Ruby Resolved

1:40PM 4/27/2017 Update: Rolling reboots are completed. 

3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address GPFS errors occured... Read more

5 years 3 months ago 5 years 3 months ago

Pages